# A Purely Data-Driven Adaptive Impedance Matching Method Robust to Parasitic Effects

### v1

@@ IV-B Strategies for Solving the Matching Problem
Ideally, accurate modeling the composite mapping $\mathcal{G}$ would enable the direct computation of the matching solution. It should be noted that the predicted $\hat{\mathbf{S}}$ from RECBM-Net may not strictly satisfy typical TMN constraints such as passivity ($\mathbf{I}-\hat{\mathbf{S}}^{\dagger}\hat{\mathbf{S}}\succeq 0$) or reciprocity ($\hat{S}_{12}=\hat{S}_{21}$). Consequently, relying solely on Eq. (16) to determine $\hat{\mathbf{S}}$ results in an underdetermined system with infinitely many solutions, meaning that the mapping $\psi^{-1}$ is multi-valued. Moreover, the circuit behavior $\mathcal{F}$ learned by RECBM-Net does not guarantee that its inverse $\mathcal{F}^{-1}$ is single-valued mapping. Above all, the composite mapping $\mathcal{G}$ may also be multi-valued mapping.
To address the challenges posed by the inherent multi-valued nature of $\mathcal{G}$ and the absence of its analytical expression, we first construct a dataset that enforces a single-valued mapping, and then train an IMS-Net to accurately approximate $\mathcal{G}$. The main process for obtaining matching solutions using IMS-Net is summarized as follows.
  • Dataset preparation: Within the TMN’s operating frequency band and tunable range, an exhaustive grid is constructed by traversing all possible combinations of frequencies and TMN configurations at a predetermined granularity. Each sample is denoted as $\mathbf{x}^{(i)}=\left\{f^{(i)},C_{p}^{(i)},C_{s}^{(i)}\right\}_{i=1}^{N_{d}}$, where $N_{d}$ is the total number of samples. Subsequently, the corresponding $\hat{\mathbf{S}}^{(i)}$ for each configuration $\mathbf{x}^{(i)}$ is predicted using the trained RECBM-Net. Assuming perfect input matching, each predicted $\hat{\mathbf{S}}^{(i)}$ and $\Gamma_{\text{in}}=0$ are substituted into Eq. (12) to compute the corresponding load reflection coefficient $\hat{\Gamma}_{L}^{(i)}$. A supervised learning dataset is then constructed as $\left\{f^{(i)},\Re(\hat{\Gamma}_{L}^{(i)}),\Im(\hat{\Gamma}_{L}^{(i)}),C_{p}^{(i)},C_{s}^{(i)}\right\}_{i=1}^{N_{d}}$, where the first three dimensions serve as input features and the last two as outputs labels. This approach not only calculates a large number of load impedances that can be matched by the practical TMN as predicted by RECBM-Net, but also directly provides the corresponding perfect matching solutions without solving the mathematical optimization problem in Eq. (15). Moreover, it guarantees a single-valued mapping from $(f,\hat{\Gamma}_{L})$ to the matching solution.
  • Offline training: IMS-Net employs the same architecture as RECBM-Net, with the only difference being the input and output dimensions (with 3 neurons for input and 2 for output). The dataset is partitioned into training and validation sets. IMS-Net is trained on the former to learn the mapping $\mathcal{G}$, and its generalization performance is evaluated on the latter.
@@
Compared with the two numerical methods discussed above, the integration of IMS-Net with RECBM-Net completely eliminates repeated RECBM-Net evaluations, thereby minimizing online computational overhead. The matching process requires only two inferences: one from RECBM-Net, to predict the the current TMN S-parameters and compute $\hat{\Gamma}_{L}$, and another from IMS-Net to directly infer the matching solution. Moreover, subsequent simulation results demonstrate that IMS-Net achieves exceptionally high accuracy. It is worth noting that IMS-Net functions as a dedicated inverse solver that is tightly coupled with its corresponding RECBM-Net. Therefore, any modification to the RECBM-Net parameters (e.g., learning different TMN circuit behaviors) necessitates retraining the associated IMS-Net.
@@ V Numerical Results and Discussion
In this section, we will first utilize RECBM-Net to characterize the circuit behavior of a practical L-network and evaluate its fitting accuracy in detail. Then, we will simulate extensive mismatch scenarios to verify the performance of the proposed data-driven adaptive impedance matching method.
@@ V-A Accuracy of RECBM-Net in Modeling Circuit Behavior
To realistically simulate the parasitic effects in practical TMN, we add numerous additional inductors, capacitors, and resistors to the ideal L-network. Fig. 6 illustrates the simulation of an L-network featuring 17 parasitic elements, with the Advanced Design System (ADS) tool employed to collect the S-parameters of the entire two-port network. We systematically sweep through all combinations of operating frequency $f$ and the values of tunable capacitors $C_{4}$ and $C_{5}$. Specifically, $f$ is varied from 1.5 GHz to 2 GHz in increments of 0.02 GHz, while $C_{4}$ and $C_{5}$ are adjusted from 0 to 10 pF in 0.02 pF steps. This comprehensive parameter sweep enabled the construction of an S-parameters dataset that accurately characterizes the practical TMN behavior, comprising a total of 6,500,000 samples.
@@
Figure 8: Comparison between predicted and ground truth S-parameters using RECBM-Net. A total of 100 samples are randomly selected from the test set containing 100,000 samples. The blue circles denote the ground truth, while the red crosses indicate the predicted values.
The dataset is divided into a training set ($80\%$) and a validation set ($20\%$). The training set is used to optimize the model parameters, while the validation set serves to evaluate the model’s generalization performance. Given the scale disparities between the frequency feature and the tunable capacitors, we apply min–max normalization to all input features to ensure training stability. In terms of the loss function, we employ the mean squared error (MSE) to quantify the discrepancy between the model’s prediction and the actual values. It is defined as
@@
where $L(\theta)$ is the loss function used to measure the average difference between the predicted vectors $\hat{\mathbf{y}}_{i}\in\mathbb{R}^{8}$ and the ground truth vectors $\mathbf{y}_{i}\in\mathbb{R}^{8}$ across all $N$ training samples. The Adam optimizer is selected to train the RECBM-Net with the initial learning rate set to $5\times 10^{-8}$. A batch size of 512 is utilized, and the model is trained for 8000 epochs.
@@
Figure 9: Empirical cumulative distribution function (ECDF) of the absolute and relative errors of RECBM-Net predictions for each real and imaginary component of the S-parameters on the test set.
@@
Figure 10: Distribution of mismatched impedances under 9000 simulated mismatch scenarios over the frequency range from 1.5 GHz to 2 GHz.
@@ V-B Performance of Data-Driven Impedance Matching Method
We simulate real-world mismatch conditions to verify the proposed purely data-driven impedance matching method and compare the performance of three matching solution determination strategies. We randomly generated 9000 combinations of $f$, $C_{4}^{*}$ and $C_{5}^{*}$ within the tunable range. Assuming the input is perfectly matched (i.e., $\Gamma_{in}=0$), the mismatched load reflection coefficient $\Gamma_{L}$ is derived based on the practical L-network topology depicted in Fig. 6. This yields 9000 sets of load impedance and corresponding TMN configurations $(C_{4}^{*},C_{5}^{*})$ that achieve perfect impedance matching. To simulate the mismatch, new values $(C_{4}^{\text{now}},C_{5}^{\text{now}})$ for tunable capacitors are generated within the tunable range and corresponding $\Gamma_{in}$ is calculated based on the practical L-network circuit model. Finally, we generated a mismatched dataset that simulates real-world impedance mismatch. Each sample comprises the operating frequency $f$, the TMN circuit configuration $(C_{4}^{\text{now}},C_{5}^{\text{now}})$, the measured input port reflection coefficient $\Gamma_{in}$, and the actual load reflection coefficient $\Gamma_{L}$ corresponding to the current mismatch scenario. Since testing the adaptive impedance matching methods within the matching forbidden region is meaningless, this simulation method ensures that the mismatched impedance $\Gamma_{L}$ remains outside matching forbidden region.
@@
Figure 11: ECDF of the absolute and relative errors between the true load reflection coefficient and one computed by applying Eq. (12) to the S-parameters predicted by RECBM-Net. The statistics are computed over 9000 simulated mismatch scenarios.
@@
Initial solution $\boldsymbol{\theta}_{0}$ (5pF, $(5\,\mathrm{pF},\,5\,\mathrm{pF})$
Learning rate $\alpha$ 0.05
Exponential decay rates $\beta_{1},\beta_{2}$ 0.9, 0.9990.90.9990.9,\,0.9990.9 , 0.999
@@
Utilizing the dataset generation method described in Sec. IV-B, we constructed the training dataset for IMS-Net. The traversal granularity for combinations of the operating frequency and TMN configurations is identical to that employed in the RECBM-Net training dataset (Sec. V-A). Similarly, we generated a supervised learning dataset formatted as $\{f,\Re(\hat{\Gamma}_{L}),\Im(\hat{\Gamma}_{L}),C_{p}^{*},C_{s}^{*}\}$ containing 6,500,000 samples with the MSE serving as the loss function. To amplify the gradient signal during backpropagation, the dataset labels were scaled by a factor of 10. Subsequently, the dataset was partitioned into $80\%$ for training and $20\%$ for validation. IMS-Net was then trained on the same hardware platform using the Adam optimizer with an initial learning rate of 0.00002. Training was conducted for 3000 epochs with a batch size of 512, ultimately converging to training and validation losses of 0.052 and 0.051, respectively. To further validate IMS-Net’s ability to predict matching solutions on unseen data, we generated a test set comprising 100,000 combinations by randomly sampling both the operating frequency and TMN configuration within a tunable range. Fig. 12 illustrates that IMS-Net reliably predicts matching solutions on the test set, with nearly $95\%$ of samples achieving a relative error below $0.6\%$. Notably, when $C_{5}^{*}$ is extremely close to 0, the prediction accuracy exhibits a minor decline, while the overall performance remains robust.
Finally, we evaluate the matching performance of the three matching solution determination strategies on the same 9,000 mismatch samples generate earlier, each reflecting a realistic mismatch scenario. Table II summarizes the hyperparameter settings for the two numerical optimization strategies. Fig. 13 illustrates that when the load reflection coefficient is derived based on the ideal L-network topology and the matching solution is calculated using the closed-form expressions in Eq. (5), the resulting solution even degrades matching performance. In contrast, the purely data-driven impedance matching method based on SAPSO achieves a reflection coefficient below 0.2 for $96.01\%$ of the mismatched samples, slightly outperforming the exhaustive grid search with a granularity of 0.01 pF $\times$ 0.01 pF, which achieves the same target for $95.76\%$ of the samples. A reflection coefficient below 0.2 is generally considered to indicate a high-quality impedance match in practical engineering applications [44], corresponding to approximately $96\%$ of the incident power being delivered to the antenna. IMS-Net also demonstrates high matching accuracy, reducing the reflection coefficient below 0.2 for $95\%$ of the samples. In comparison, the AD-Adam method yields lower matching accuracy, with $88.95\%$ of the samples satisfying the same reflection coefficient threshold. This suboptimal performance can be attributed to the inherent limitations of gradient-based optimization algorithms when applied to non-convex objective landscapes.
@@
Figure 13: Comparison of the matching performance of solutions obtained by the proposed data-driven approach with different determination strategies, those analytically derived from the ideal L-network circuit model.
@@
Figure 14: ECDF of the number of DNN inferences required by three matching solution determination strategies (SAPSO, AD-Adam, and IMS-Net) across 9,000 simulated mismatch scenarios.
@@
Fig. 14 presents the online computational overhead of the three matching solution computation strategies, evaluated over 9,000 simulated mismatch scenarios. The overhead is quantified by the number of DNN inferences required, and their distributions are compared using ECDFs. SAPSO achieves the highest matching accuracy among the three strategies, but incurs the highest computational overhead. It requires thousands of RECBM-Net inferences per sample due to its population-based search process. In contrast, AD-Adam typically requires dozens of DNN inferences. Furthermore, IMS-Net computes the matching solution using only a single DNN inference, making it highly efficient for real-time applications. Although AD-Adam yields lower matching accuracy, it offers a substantial advantage in computational efficiency compared to SAPSO. IMS-Net achieves both high matching accuracy and minimal online computational overhead but it requires an offline training phase, which introduces additional computational overhead prior to deployment. In summary, the proposed purely data-driven adaptive impedance matching method does not rely on explicit circuit topology and remain effective and accurate when the practical circuit topology deviates significantly from the ideal TMN.

### v2

@@ IV-B Strategies for Solving the Matching Problem
Ideally, accurately modeling the composite mapping $\mathcal{G}$ would enable the direct computation of the matching solution. It should be noted that the predicted $\hat{\mathbf{S}}$ from RECBM-Net may not strictly satisfy typical TMN constraints such as passivity ($\mathbf{I}-\hat{\mathbf{S}}^{\dagger}\hat{\mathbf{S}}\succeq 0$) or reciprocity ($\hat{S}_{12}=\hat{S}_{21}$). Consequently, relying solely on Eq. (16) to determine $\hat{\mathbf{S}}$ results in an underdetermined system with infinitely many solutions, meaning that the mapping $\psi^{-1}$ is multi-valued. Moreover, the circuit behavior $\hat{\mathcal{F}}$ learned by RECBM-Net does not guarantee that its inverse $\hat{\mathcal{F}}^{-1}$ is a single-valued mapping. Above all, the composite mapping $\mathcal{G}$ may also be multi-valued mapping.
To address the challenges posed by the inherent multi-valued nature of $\mathcal{G}$ and the absence of its analytical expression, we first construct a dataset that enforces a single-valued mapping, and then train an IMS-Net to accurately approximate $\mathcal{G}$. The main process for obtaining matching solution using IMS-Net is summarized as follows.
  • Dataset preparation: Within the TMN’s operating frequency band and tunable range, an exhaustive grid is constructed by traversing all possible combinations of frequencies and TMN configurations at a predetermined granularity. Each sample is denoted as $\mathbf{x}^{(i)}=\left\{f^{(i)},C_{p}^{(i)},C_{s}^{(i)}\right\}_{i=1}^{N_{d}}$, where $N_{d}$ is the total number of samples. Subsequently, the corresponding $\hat{\mathbf{S}}^{(i)}$ for each configuration $\mathbf{x}^{(i)}$ is predicted using the trained RECBM-Net. Assuming perfect input matching, each predicted $\hat{\mathbf{S}}^{(i)}$ and $\Gamma_{\text{in}}=0$ are substituted into Eq. (12) to compute the corresponding load reflection coefficient $\hat{\Gamma}_{L}^{(i)}$. A supervised learning dataset is then constructed as $\left\{f^{(i)},\Re(\hat{\Gamma}_{L}^{(i)}),\Im(\hat{\Gamma}_{L}^{(i)}),C_{p}^{(i)},C_{s}^{(i)}\right\}_{i=1}^{N_{d}}$, where the first three dimensions serve as input features and the last two as output labels. This approach not only calculates a large number of load impedances that can be matched by the practical TMN as predicted by RECBM-Net, but also directly provides the corresponding perfect matching solutions without solving the mathematical optimization problem in Eq. (15). Moreover, it guarantees a single-valued mapping from $(f,\hat{\Gamma}_{L})$ to the matching solution.
  • Offline training: IMS-Net employs the same architecture as RECBM-Net, with the only difference being in the output dimension (two neurons in the output layer). The dataset is partitioned into training and validation sets. IMS-Net is trained on the former to learn the mapping $\mathcal{G}$, and its generalization performance is evaluated on the latter.
@@
Compared with the two numerical methods discussed above, the integration of IMS-Net with RECBM-Net completely eliminates repeated RECBM-Net evaluations, thereby minimizing online computational overhead. The matching process requires only two inferences: one from RECBM-Net, to predict the current TMN S-parameters and compute $\hat{\Gamma}_{L}$, and another from IMS-Net to directly infer the matching solution. Moreover, subsequent simulation results demonstrate that IMS-Net achieves exceptionally high accuracy. It is worth noting that IMS-Net functions as a dedicated inverse solver that is tightly coupled with its corresponding RECBM-Net. Therefore, any modification to the RECBM-Net parameters (e.g., learning different TMN circuit behaviors) necessitates retraining the associated IMS-Net.
@@ V Numerical Results and Discussion
In this section, we will first utilize RECBM-Net to characterize the circuit behavior of a practical L-network and evaluate its fitting accuracy in detail. Then, we will simulate extensive mismatched scenarios to verify the performance of the proposed data-driven adaptive impedance matching method.
All algorithm development and testing are carried out in a Python environment (version 3.8.20). The hardware platform is a workstation equipped with an Intel Xeon Gold 5218 central processing unit (CPU) @ 2.30 GHz and four NVIDIA GeForce RTX 2080 Ti graphics processing units (GPUs). The analytical method for computing the ABCD matrix to indirectly derive the S-parameters, as described in Section III-A, is implemented in Python and verified to be consistent with simulation results obtained from Keysight Advanced Design System. This implementation is employed to generate the training and testing datasets for RECBM-Net by computing S-parameters across various circuit configurations. Similarly, the training and testing datasets for IMS-Net, as well as the simulated mismatched scenarios used for performance evaluation, are all generated in Python. The detailed data generation procedures are provided in Sections IV-B and V-B. Additionally, both RECBM-Net and IMS-Net are trained using the PyTorch framework (version 2.4.1). The SAPSO and AD-Adam are further refined through hyperparameter tuning to improve performance. Finally, the proposed matching solution determination strategies are implemented in Python and employed to validate their end-to-end matching performance.
@@ V-A Accuracy of RECBM-Net in Modeling Circuit Behavior
To simulate the deviation of circuit behavior from the ideal matching network caused by practical parasitic effects, we introduce additional parasitic elements into the ideal L-network structure. Further, ladder circuits with alternating series and parallel branches are incorporated to simulate additional parasitic effects. Fig. 6 illustrates the simulation of an L-network featuring 17 parasitic elements. The ladder topology is adopted because it is widely used in engineering to approximate the system response of practical devices. According to vector fitting techniques [39] and network synthesis theory [40], the system response of a practical device under test (DUT) can be approximated by rational functions, which can subsequently be synthesized into a ladder-type equivalent circuit. Therefore, based on this structure, we randomly introduce a large number of additional inductors, capacitors, and resistors, with their values also assigned randomly. The objective is to construct a high-order, sufficiently complex test circuit to validate our data-driven impedance matching method. In this circuit, analytical matching solutions derived from the ideal L-network model fail completely. In our proposed method, the internal circuit structure is treated as unknown, in order to verify its effectiveness when parasitic effects obscure the practical configuration.
We generate the S-parameter dataset for the practical L-network using the Python script that analytically computes the S-parameters of a given two-port network. We systematically sweep through all combinations of operating frequency $f$ and the values of tunable capacitors $C_{4}$ and $C_{5}$. Specifically, $f$ is varied from 1.5 GHz to 2 GHz in increments of 0.02 GHz, while $C_{4}$ and $C_{5}$ are adjusted from 0 to 10 pF in 0.02 pF steps. This comprehensive parameter sweep enables the construction of an S-parameter dataset that accurately characterizes the practical TMN behavior, comprising a total of 6,500,000 samples. The generation of this dataset takes 202.28 seconds on an Intel Xeon Gold 5218 CPU.
@@
Figure 8: Comparison between predicted and ground truth S-parameters using RECBM-Net. A total of 50 samples are randomly selected from the test set containing 100,000 samples. The blue circles denote the ground truth, while the red crosses indicate the predicted values.
The dataset is divided into a training set ($80\%$) and a validation set ($20\%$). The training set is used to optimize the model parameters, while the validation set serves to evaluate the model’s generalization performance. Since the input features include frequency and tunable capacitor values with different numerical scales, we apply min–max normalization to all features to ensure training stability. In terms of the loss function, we employ the mean squared error (MSE) to quantify the discrepancy between the model’s prediction and the actual values. It is defined as
@@
where $L(\theta)$ measures the average difference between the predicted vectors $\hat{\mathbf{y}}_{i}\in\mathbb{R}^{8}$ and the ground truth vectors $\mathbf{y}_{i}\in\mathbb{R}^{8}$ across all $N$ training samples. The Adam optimizer is selected to train the RECBM-Net with an initial learning rate set to $5\times 10^{-8}$. A batch size of 512 is utilized, and the model is trained for 8000 epochs. Training is performed with PyTorch’s DistributedDataParallel (DDP) framework on four NVIDIA GeForce RTX 2080 Ti GPUs, requiring approximately 59.39 hours in total.
@@
Figure 9: ECDF of the absolute and relative errors of RECBM-Net predictions for each real and imaginary component of the S-parameters on the test set.
@@
Figure 10: Distribution of mismatched impedances under 9000 simulated mismatched scenarios over the frequency range from 1.5 GHz to 2 GHz.
@@ V-B Performance of Data-Driven Impedance Matching Method
We simulate real-world mismatch conditions to verify the proposed data-driven impedance matching method and compare the performance of three matching solution determination strategies. We randomly generate 9000 combinations of $f$, $C_{4}^{*}$ and $C_{5}^{*}$ within the tunable range. Assuming the input is perfectly matched (i.e., $\Gamma_{in}=0$), the mismatched load reflection coefficient $\Gamma_{L}$ is derived based on the practical L-network topology depicted in Fig. 6. This yields 9000 sets of load impedance and corresponding TMN configurations $(C_{4}^{*},C_{5}^{*})$ that achieve perfect impedance matching. To simulate the mismatch, new values $(C_{4}^{\text{now}},C_{5}^{\text{now}})$ for tunable capacitors are generated within the tunable range and the corresponding $\Gamma_{\text{in}}$ is calculated based on the practical L-network circuit model. Finally, we generate a mismatched dataset that simulates real-world impedance mismatch. Each sample comprises the operating frequency $f$, the TMN circuit configuration $(C_{4}^{\text{now}},C_{5}^{\text{now}})$, the measured input port reflection coefficient $\Gamma_{\text{in}}$, and the actual load reflection coefficient $\Gamma_{L}$ corresponding to the current mismatched scenario. Since testing the adaptive impedance matching methods within the matching forbidden region is meaningless, this simulation method ensures that the mismatched impedance $\Gamma_{L}$ remains outside the matching forbidden region.
@@
Figure 11: ECDF of the absolute and relative errors between the true load reflection coefficient and that calculated from Eq. (12) using the S-parameters predicted by RECBM-Net. The statistics are computed over 9000 simulated mismatched scenarios.
@@
Initial solution $\boldsymbol{\theta}_{0}$ $(5\,\mathrm{pF},\,5\,\mathrm{pF})$
Learning rate $\alpha$ 0.013
Exponential decay rates $\beta_{1},\beta_{2}$ 0.9, 0.9990.9,\,0.999
@@
Utilizing the dataset generation method described in Section IV-B, we construct the training dataset for IMS-Net. The traversal granularity for combinations of the operating frequency and TMN configurations is identical to that employed in the RECBM-Net training dataset (Section V-A). Similarly, we generate a supervised learning dataset formatted as $\{f,\Re(\hat{\Gamma}_{L}),\Im(\hat{\Gamma}_{L}),C_{p}^{*},C_{s}^{*}\}$ containing 6,500,000 samples with the MSE serving as the loss function. To amplify the gradient signal during backpropagation, the dataset labels (in pF) are scaled by a factor of 10. Subsequently, the dataset is partitioned into $80\%$ for training and $20\%$ for validation. IMS-Net is then trained on the same hardware platform using the Adam optimizer with an initial learning rate of 0.00002. Training is conducted for 3000 epochs with a batch size of 512, ultimately converging to training and validation losses of 0.052 and 0.051, respectively. To further validate IMS-Net’s ability to predict matching solutions on unseen data, we generate a test set comprising 100,000 combinations by randomly sampling both the operating frequency and TMN configuration within a tunable range. Fig. 12 illustrates that IMS-Net reliably predicts matching solutions on the test set, with nearly $95\%$ of samples achieving a relative error below $0.6\%$. Notably, when $C_{5}^{*}$ is extremely close to 0, the prediction accuracy exhibits a minor decline, while the overall performance remains robust.
Finally, we evaluate the end-to-end performance of the three matching solution determination strategies on the same 9000 mismatched samples generated earlier. For a fair comparison and improved computational efficiency, an early stopping threshold $\varepsilon$ of 0.005 is applied to both methods. Empirical observations confirm that this threshold is reliably attainable and corresponds to excellent matching performance. For SAPSO, following the parameter settings in [41], both the individual learning factor $\kappa_{1}$ and the social learning factor $\kappa_{2}$ are set to 2.05, while the cooling factor $\lambda$ is set to 0.5. The number of particles $N$ and the maximum iterations $\mathcal{I}_{\max}$ are determined through hyperparameter tuning. For AD-Adam, the initial solution is set to the midpoint of the tunable capacitance range to reduce the number of required iterations. The exponential decay rates $\beta_{1}=0.9$, $\beta_{2}=0.999$, together with the stability constant $\epsilon=10^{-8}$ are set according to the configuration in [37]. The learning rate $\alpha$ and the maximum iterations $\mathcal{I}_{\max}$ are also determined through hyperparameter tuning. The detailed hyperparameter tuning procedures for both SAPSO and AD-Adam are provided in Appendix B of Supplementary Material. Table II summarizes the hyperparameter settings for the two numerical optimization strategies.
[…]
(a) ECDF of median.
(b) ECDF of SD.
Figure 13: ECDFs of the median and standard deviation of the reflection coefficient magnitude obtained by SAPSO over 30 independent runs per mismatched scenario across 9000 scenarios.
[…]
In addition to the ECDFs of the tuned reflection coefficient magnitude for each matching strategy, we also summarize the overall mean, median, and SD of the tuned reflection coefficient magnitudes across all mismatched scenarios. As shown in Table III, both SAPSO and IMS-Net achieve mean and median reflection coefficient magnitudes well below 0.2, with standard deviations around 0.1. These descriptive statistics demonstrate that both methods can consistently achieve low reflection coefficient magnitudes. Although AD-Adam exhibits a relatively larger SD, its mean and median remain well below 0.2. These results indicate that, despite its higher variability, the AD-Adam method is still capable of achieving effective impedance matching in most mismatched scenarios, even in the presence of parasitic effects.
@@
Figure 14: Comparison of the matching performance of solutions obtained by the proposed data-driven method with different determination strategies, those analytically derived from the ideal L-network circuit model.
@@
Figure 15: ECDF of online computational overhead for three matching solution determination strategies (SAPSO, AD-Adam, and IMS-Net) across 9000 simulated mismatched scenarios. The computational overhead is quantified by the equivalent number of DNN inferences required.
@@
Fig. 15 presents the online computational overhead of the three matching solution determination strategies, evaluated over 9000 simulated mismatched scenarios. The overhead is quantified by the number of DNN inferences required, and their distributions are compared using ECDFs. Each strategy requires one RECBM-Net inference to compute the load reflection coefficient, which is included in the analysis. SAPSO achieves the highest matching accuracy among the three strategies, but incurs the highest computational overhead. Owing to its population-based search process, each sample requires an average of 2097 RECBM-Net inferences. For the AD-Adam method, the computational overhead of gradient evaluation in each iteration is approximately equivalent to that of a single RECBM-Net inference. Consequently, determining the matching solution with AD-Adam requires an average of 285 DNN inferences per sample. Furthermore, IMS-Net requires only one additional DNN inference for the matching solution, resulting in two inferences in total when including the inference for the load reflection coefficient. Although AD-Adam yields lower matching accuracy, it offers a substantial advantage in computational efficiency compared to SAPSO. IMS-Net achieves both high matching accuracy and minimal online computational overhead, but it requires an offline training, which introduces additional computational overhead prior to deployment. In summary, the proposed data-driven impedance matching method does not rely on explicit circuit topology and remains effective and accurate when the practical circuit topology deviates significantly from the ideal TMN.
TABLE III: Descriptive statistics of the tuned reflection coefficient magnitudes for the three matching solution determination strategies
Matching strategy Mean Median SD
RECBM-Net+Grid Search 0.03062 0.00314 0.10214
RECBM-Net+SAPSO 0.03097 0.00403 0.10099
RECBM-Net+AD-Adam 0.06531 0.00648 0.20491
RECBM-Net+IMS-Net 0.03717 0.00323 0.12107
Ideal L-network 0.90261 0.97546 0.16649
In addition to the performance analysis of the proposed data-driven impedance matching method, we also report the overall turnaround time for the algorithm development and testing using Python. First, RECBM-Net is trained to learn the circuit behavior. The entire process of dataset generation, model training, and testing takes approximately 59.45 hours. Next, to improve the performance of the numerical optimization methods, we derive a penalty function for SAPSO and perform hyperparameter tuning for both the SAPSO and AD-Adam methods. The corresponding development times are 39.54, 16.95, and 3.93 hours, respectively. Subsequently, IMS-Net is developed to directly predict matching solutions, where dataset generation, model training, and testing collectively require approximately 20.05 hours. Finally, to evaluate the end-to-end performance of the three matching solution determination strategies, we simulate 9000 mismatched scenarios. Dataset generation requires only 0.1422 seconds, while the evaluation times for SAPSO, AD-Adam, and IMS-Net across the 9000 scenarios are approximately 106.5 hours, 2.37 hours, and 0.00144 seconds, respectively. The development process described above is implemented sequentially on a single workstation. It is worth noting that the SAPSO development involves repeated experiments, which could be parallelized across multiple workstations to reduce the overall runtime. In summary, the overall experimental turnaround time for the proposed data-driven impedance matching method is approximately 248.79 hours. This time can be further reduced through the use of additional computing resources, highlighting the method’s potential to shorten the time-to-market (TTM).
V-C Performance under imperfect measurements
In this section, we first examine the impact of parasitic effects in the DUT on reflection coefficient measurement approaches applicable to small and portable platforms. A six-port reflectometer [43] provides a simple yet accurate means of measuring the reflection coefficient. It consists of a passive six-port network and four power detectors. The reflection coefficient of the DUT is computed from the scalar power measurements at the four detector ports. Since the computation depends solely on measured power and requires no knowledge of the DUT’s internal circuit parameters, the resulting reflection coefficient inherently captures the effects of DUT parasitics rather than being distorted by them. Similarly, the three-point measurement method [44, 17] based on a uniform transmission line offers another low-cost and practical solution for reflection coefficient measurement. In this approach, the DUT’s reflection coefficient is computed from voltage magnitudes sampled at three distinct locations along the line. This computation is likewise independent of the DUT’s internal characteristics, ensuring that the measured reflection coefficient is not biased by parasitic effects.
[…]
Imperfect measurement of impedance mismatch metrics is a common issue that leads to degraded performance in adaptive impedance matching methods. The primary focus of this paper, however, is on developing an efficient and accurate adaptive impedance matching method that remains effective in the presence of parasitic effects.
V-D Feasibility analysis of embedded platform deployment
To assess the feasibility of implementing the proposed deep learning-based impedance matching methods in the radio unit illustrated in Fig. 1, we analyze their computational resource requirements. Specifically, we measure the static memory usage during deployment, the theoretical floating-point operations (FLOPs) associated with the deep learning component of the optimization process, and the peak runtime memory usage of each method.
RECBM-Net and IMS-Net contain 1,396,296 and 1,395,906 parameters, respectively, each requiring approximately 5.58 megabytes (MB) of single-precision storage. The number of parameters involved in the solution update procedures of SAPSO and AD-Adam is negligible compared with those of the DNN models. Consequently, both SAPSO and AD-Adam implementations require approximately 5.58 MB of static memory, whereas the method that directly determines the matching solution using IMS-Net requires $5.58\times 2=11.16$ MB. Across all three methods, the primary computational bottleneck arises from deep learning operations. Therefore, we measure the theoretical computational overhead of a single inference for RECBM-Net and IMS-Net to be 2.7872 MFLOPs and 2.7864 MFLOPs, respectively. Across the tested mismatched samples, the SAPSO method requires an average of 2097 RECBM-Net inferences per sample. Accounting for gradient computation overhead, the AD-Adam method incurs an average computational overhead equivalent to 285 RECBM-Net inferences for each mismatched sample. By contrast, IMS-Net obtains the matching solution with only one inference from RECBM-Net and one from IMS-Net. Multiplying the number of inferences by the corresponding theoretical inference FLOPs yields estimated computational overheads of 5.8448 GFLOPs, 0.7943 GFLOPs, and 0.0055 GFLOPs for the SAPSO, AD-Adam, and IMS-Net methods, respectively. Finally, the measured peak runtime memory usages for continuously determining matching solutions across mismatched samples are 587.35 MB for SAPSO, 792.62 MB for AD-Adam, and 362.09 MB for IMS-Net. A summary of these results is provided in Table IV. Although these metrics are obtained on an Intel Xeon Gold 5218 CPU, they characterize the inherent computational resource demands of the methods and thus provide a basis for evaluating deployment feasibility on embedded systems.
[…]
In summary, the proposed deep learning-based impedance matching methods can theoretically be deployed on two representative embedded platforms. On hardware with lower computational capability, the SAPSO method may fall short of meeting real-time requirements. In contrast, the AD-Adam and IMS-Net methods can achieve millisecond-level real-time matching. The IMS-Net method, in particular, exhibits negligible latency in determining the matching solution, making it well suited for impedance matching applications that demand strict real-time performance.
TABLE IV: Computational complexity metrics for the three methods.
Metric SAPSO AD-Adam IMS-Net
Static storage size (MB) 5.58 5.58 11.16
Theoretical FLOPs (GFLOPs) 5.8448 0.7943 0.0055
Peak memory usage (MB) 587.35 792.62 362.09